In the past, patrons of a cinema complex have spent an average of $5.00 for popcorn and other snacks, with a standard deviation of $1.80. The amounts of these expenditures have been normally distributed. Following an intensive publicity campaign by a local medical society, the mean expenditure for a sample of 18 patrons is found to be $4.20. In a one-tail test at the 0.05 level of significance, does this recent experience suggest a decline in spending? (Critical Value Approach)
Answer #1

x̅ = 4.2, σ = 1.8, n = 18

Null and Alternative hypothesis:

Ho : µ = 5 ; H1 : µ < 5

Critical value, z crit = NORM.S.INV(0.05) = -1.645

Reject Ho if z < -1.645

Test statistic:

z = (x̅- µ)/(σ/√n) = (4.2 - 5)/(1.8/√18) = -1.8856

Decision:

As z = -1.8856 > -1.645, Reject the null hypothesis

Conclusion:

There is enough evidence to conclude that there is decline in spending at 0.05 significance level.